Translate the decimal number 167 into binary and hexadecimal. Show your work.
Answer #1

1)
Divide 167 successively by 2 until the quotient is 0
   > 167/2 = 83, remainder is 1
   > 83/2 = 41, remainder is 1
   > 41/2 = 20, remainder is 1
   > 20/2 = 10, remainder is 0
   > 10/2 = 5, remainder is 0
   > 5/2 = 2, remainder is 1
   > 2/2 = 1, remainder is 0
   > 1/2 = 0, remainder is 1
Read remainders from the bottom to top as 10100111
So, 167 of decimal is 10100111 in binary
Answer: 10100111

2)
Divide 167 successively by 16 until the quotient is 0
167/16 = 10, remainder is 7
10/16 = 0, remainder is 10
Read remainders from the bottom to top as A7
Answer: 0xA7